# ALMA FACTS III. High-Resolution CO(2-1)/CO(1-0) Maps of Twelve Nearby Galaxies

### v1

@@ Abstract
We present early results from a high-resolution analysis (100-200pc) of the CO(2-1)/CO(1-0) line ratio in twelve nearby galaxies. We use new ALMA CO(1-0) observations from the Fundamental CO(1-0) Transition Survey (FACTS), and re-imaged CO(2-1) data from PHANGS. We make empirical classifications based on the optical and molecular gas morphologies, which show clear systematic trends in the variation of as a function of galactic structure. The sample includes barred and unbarred, and flocculent galaxies. The barred spiral galaxies follow a general trend: is high in the center, low along the bar, increases at the bar ends, and then declines in the outer parts of the disk. The structure dependence suggests the importance of galactic dynamics on molecular gas evolution, and consequently on star formation, in galaxies. fluctuates in the spiral arms for both barred and unbarred galaxies. HII regions increase locally in their surrounding gas and are often associated with galactic structures. Together, varies systematically as a function of galactic structure, dynamics, and star formation activity.
1 Introduction
@@
Studies in the Milky Way (MW) have shown systematic variations in within and amongst molecular clouds, and in the Galactic disk (Hasegawa, 1997, for a review). These studies have classified molecular gas according to as low-ratio gas (LRG) with , high-ratio gas (HRG) with , and very-high ratio gas (VHRG) . For example, varies between and within star-forming and non-star-forming molecular clouds. Sakamoto et al. (1994) and more recently Nishimura et al. (2015) mapped the star-forming Orion molecular clouds, and found that ranges from 0.5 to 1.0 between the main ridge and peripheries of Orion A. Falgarone et al. (1998) studied non-star-forming molecular clouds and found they have roughly uniform 0.65. In the Galactic disk, radially decreases (Sakamoto et al., 1997), with azimuthal variations, i.e., LRG in the interarms and HRG in the spiral arms. In the Galactic center, the molecular gas exhibits distinctive properties than gas in the disk, and can be excited to high , and occasionally, very-high ratios (Oka et al., 1996, 1998b; Sawada et al., 2001). Overall, the ratio varies from 0.96 in the central kiloparsec region of the MW, while the typical disk value is 0.6-0.7 (Sawada et al., 2001). VHRG with is relatively rare and is typically found in the close vicinity of star formation, e.g. due to interaction with supernovae, or in the galactic center (Oka et al., 1996; Seta et al., 1998).
@@
There have been many efforts to characterize in external galaxies. However, previous measurements were often hindered by challenges in calibration, especially since intrinsically has a relatively small dynamic range (only a factor of 2-3 variation; Garcia-Burillo et al., 1993; Crosthwaite et al., 2002; Lundgren et al., 2004; Crosthwaite & Turner, 2007; Koda et al., 2012; Yajima et al., 2021; Leroy et al., 2022). The calibration issue often led to high over large regions of the galaxies, seemingly indicating widespread optically-thin CO gas, which is not very likely. Spatial variations in in external galaxies began to become apparent with improvements in calibration and techniques (e.g., Koda et al., 2012). In particular, with the advent of ALMA, studies of have become increasingly possible (Koda et al., 2020; Egusa et al., 2022; Maeda et al., 2022; den Brok et al., 2023).
@@
High-resolution studies of remain relatively scarce and are typically case-studies, or due to limited coverage, focus on particular regions in galaxies (Egusa et al., 2022; Maeda et al., 2022; den Brok et al., 2023; Koda et al., 2025). These studies have found values similar to those of Galactic studies as well as indications of trends with respect to galactic structures. Beyond case-studies, it is clear that a survey of in nearby galaxies at high-resolution and with well-calibrated data, is needed to understand the general evolution of molecular gas and their physical conditions. While there are existing surveys of variations (e.g., Yajima et al., 2021; Leroy et al., 2022), they are at lower sensitivity and resolution, and the compared CO(1-0) and CO(2-1) data were often observed from different telescopes. As a result, overall systematic trends were not clear.
@@
In this paper, we analyze in twelve nearby galaxies using new high-resolution CO(1-0) data from the ALMA-FACTS survey and reimaged CO(2-1) data from the PHANGS-ALMA survey (Leroy et al., 2021). We present early results, which already empirically show systematic trends as a function of galactic structure. This is the third paper of the ALMA-FACTS Survey. on kpc-scales in the same twelve sample galaxies using the ALMA Total Power (TP) component has been studied (ALMA-FACTS Paper II; Komugi et al., 2025). We plan multi-wavelength comparisons for future studies.
2 Observations and Data Reduction
@@
This section discusses the data reduction, mainly on the CO(2-1) data (Section 2.1). The CO(1-0) data reduction is discussed in a separate paper (Koda et al. in prep.), but is practically the same with one note in Section 2.2. This method has been discussed in depth and used by Koda et al. (2023) and Koda et al. (2025).
@@
Figure 1 shows the sample in Spitzer 3.6 m dust emission from S4G (Sheth et al., 2010; S4G Team, 2020) and SINGS (for NGC 3621, Kennicutt et al., 2003; SINGS Team, 2020) to demonstrate the field-of-view (FoV) of the observations. The FoVs of the 12m+07m CO(2-1) observations (yellow contours), and consequently the maps (section 4.1), cover the inner disks of the galaxies.
2.1 CO(2-1)
@@ 2.1.1 Calibration
2.1.2 Imaging
The 12m and 07m array data for each galaxy were imaged together in MIRIAD, instead of CASA, because it takes into account spatial variations in the point spread function (PSF) (Koda et al., 2023). To prepare for imaging in MIRIAD, we used the CASA tasks split and mstransform to split the narrowband spectral window (SPW) containing the CO(2-1) emission out from each 12m and 07m MS, and then to regrid them to a common velocity grid with channel width 5 km/s. These were converted into MIRIAD format. We used the average system temperature Tsys and integration times tint from each MS following Koda et al. (2023). The invert task in MIRIAD was then performed on these 12m and 07m visibilities to produce the CO(2-1) data cubes. We used robust = +2 (corresponding to natural weighting), cellsize = 0.25”, and a velocity coverage and width of matching those of the CO(1-0) data cubes (Koda et al. in prep).
@@ 2.1.2 Imaging
2.2 CO(1-0)
The data reduction for CO(1-0) is discussed in a separate paper (Koda et al., in prep.), but is practically the same as the one discussed above. The only difference is the very first step of imaging, where we did not have a reference cube to make a mask. We went through the full imaging process without any mask and generated a 12m+07m+TP cube in CO(1-0). A mask is made with this cube for the second imaging as discussed above.
3 Maps
@@
To make the integrated intensity maps for each galaxy, we smoothed the CO(1-0) and CO(2-1) cubes to a common angular resolution of 2.5. We then regridded the CO(2-1) data spatially to the CO(1-0) data and applied a mask to the data cubes. The mask was based on the (cleaned) CO(1-0) signal-to-noise (SN) cube starting from volumes that have at least 80 pixels with 5.0. This mask was then expanded spatially by 2.5 (=10 pixels in length) to include potential diffuse extended emission (if it exists). After applying this mask, the cubes were integrated along the velocity axis to produce the integrated intensity maps, and corresponding error maps were generated. The error in each pixel is calculated as RMS, where RMS is the RMS noise per channel width , and is the number of channels included in the velocity integration.
These are shown in Figures 3 and 4. The colorbar scales for are adjusted to cover 0.7 times the shown range of for a fairer comparison using a (fiducial) average line ratio of 0.7. The and maps appear similar for prominent structures, although their ratio maps show variations (see Section 4).
@@
(7-9) Adopted semi-major axis, position-angle, and ellipticity of the bar from Herrera-Endoqui et al. (2015) for bar region masks (section 5.2.5).
@@
4 Results
From the integrated intensity maps in Section 3, we calculated the line ratio maps (Figures 5-10). We applied a signal-to-noise (SN) cut of 5 in CO(1-0) to ensure enough significance in the denominator of the ratio. While the CO(1-0) sensitivity is similar among all the galaxies, the CO(2-1) sensitivity varies between galaxies. Hence, we impose a less stringent cut of 3 in CO(2-1) to mitigate the effect of noise.
We summarize the values and their distribution for each galaxy individually. The spatial resolutions of our maps at the adopted angular resolution of 2.5 range from 50-200 pc amongst the sample (Table 1), larger than the typical size of a molecular cloud. The values presented here thus represent the average within individual molecular clouds.
4.1 Maps and Distributions
The maps for each galaxy and their distributions (histograms), as well as their optical images for comparison, are shown in Figures 5-10. We summarize the variations in each galaxy separately (see Table 3). We tend to find similar trends in each morphological type and galactic structure (e.g., barred, unbarred, spiral), which leads us to classify them according to their optical and molecular gas morphologies (Section 5).
We divide the molecular gas using the definitions adopted in Hasegawa (1997, for a review). Low-ratio gas (LRG) have (blue), high-ratio gas (HRG) are in between (red), and very-high-ratio gas (VHRG) have . The division of molecular gas at (white) is motivated both theoretically via the LVG approximation (Goldreich & Kwan, 1974; Scoville & Solomon, 1974) and observationally first in MW studies (e.g., Sakamoto et al., 1997; Oka et al., 1998a; Hasegawa, 1997). Typical molecular gas with physical conditions () = (300 cm-3, 10 K) have 0.6-0.7 (i.e., LRG), with the transition from LRG to HRG requiring 2-3x denser and/or hotter gas (e.g., Koda et al., 2012). On the other hand, VHRG with is difficult to achieve, requiring optically thin conditions yet even denser/hotter gas.
@@ 4.1 Maps and Distributions
4.1.1 NGC 0628
NGC 0628 is an isolated spiral galaxy (SAc) with many star-forming regions (HII) along its spiral arms. It does not have a bar structure. At its distance, the angular resolution 2.5 corresponds to a spatial resolution of 119 pc. From Figure 3, the molecular gas distribution in NGC 0628 is relatively dim in both transitions of CO and its spiral arms appear loosely wound. The stellar spiral arms extend beyond the CO coverage (e.g., Figure 1).
The measurements are concentrated around the spiral arms due to the sensitivity limitation, with that in the interarm regions only occasionally determined. The disk is dominated by low-ratio gas (Figure 5), encompassing 70% of the significantly detected pixels, while high-ratio gas makes up 24% (blue histogram). This is similar for the CO(1-0) intensity distribution (green histogram), with low-ratio gas and high-ratio gas containing 78% and 18% respectively. Over the whole area, the mean is 0.62 while the intensity-weighted mean is 0.59.
4.1.2 NGC 1097
@@
is high in the center, becomes lower along the bar, and then increases at the bar ends (Figure 5). Low-ratio gas and high-ratio gas make up 73% and 24% of the area (blue histogram), respectively. However, the CO(1-0) intensity contained by low-ratio and high-ratio gas are similar, at 48% and 50% respectively (green histogram). Over the whole area, the mean is 0.61 while the intensity-weighted mean is 0.70.
@@ 4.1.3 NGC 1512
is high in the center and becomes lower along the offset ridges (Figure 6). The along the spiral arms fluctuate between low-ratio and high-ratio, with some high-ratio gas spatially coincident with HII regions. This galaxy also has a central gas concentration and high ratio region larger than the nominal definition of the central 1 kpc. Indeed, the histogram for the central 2 kpc explains most of the HRG. Low-ratio and high-ratio gas encompass 58% and 36% of the area shown, respectively. However, low-ratio gas contain 37% of the CO(1-0) intensity, while 56% is contained within the high-ratio gas. Over the whole area, the mean is 0.67 while the intensity-weighted mean is 0.76. From the histograms, the central 1-2 kpc ( kpc) contain mostly high-ratio gas.
4.1.4 NGC 3351
@@
In Figure 6, is high in the center and decreases along the offset ridges, but fluctuates in the surrounding spiral arms between low and high ratios. Low-ratio and high-ratio gas make up 67% and 30% of the area shown (blue histogram). However, the amount of low-ratio gas and high-ratio gas contained by the CO(1-0) intensity are similar at 44% and 55%, respectively (green histogram). Over the whole area, the mean is 0.65 while the intensity-weighted mean is 0.73. From the histograms, high-ratio gas is concentrated predominantly in the central kiloparsec ( kpc) region.
4.1.5 NGC 3521

### v2

@@ Abstract
We present early results from a high-resolution analysis (100-200pc) of the CO(2-1)/CO(1-0) line ratio in twelve nearby galaxies. We use new ALMA CO(1-0) observations from the Fundamental CO(1-0) Transition Survey (FACTS), and re-imaged CO(2-1) data from PHANGS. We make empirical classifications based on the optical and molecular gas morphologies, which show clear systematic trends in the variation of as a function of galactic structure. The sample includes barred and unbarred, and flocculent galaxies. The barred spiral galaxies follow a general trend when the gas exists significantly: is high in the center, low along the bar, increases at the bar ends, and then lowers beyond the bar end or flattens in the outer parts of the disk. The structure dependence suggests the importance of galactic dynamics on molecular gas evolution, and consequently on star formation, in galaxies. fluctuates in the spiral arms for both barred and unbarred galaxies. Areas around HII regions in some cases appear to show more high-ratio gas. Together, varies systematically as a function of galactic structure, dynamics, and star formation activity.
I Introduction
@@
Studies in the Milky Way (MW) have shown systematic variations in within and amongst molecular clouds, and in the Galactic disk (see Hasegawa, 1997). These studies have classified molecular gas according to as low-ratio gas (LRG) with , high-ratio gas (HRG) with , and very-high ratio gas (VHRG) . For example, varies between and within star-forming and non-star-forming molecular clouds. Sakamoto et al. (1994) and more recently Nishimura et al. (2015) mapped the star-forming Orion molecular clouds, and found that ranges from 0.5 to 1.0 between the main ridge and peripheries of Orion A. Falgarone et al. (1998) studied non-star-forming molecular clouds and found they have roughly uniform 0.65. In the Galactic disk, radially decreases (Sakamoto et al., 1997), with azimuthal variations, i.e., LRG in the interarms and HRG in the spiral arms. In the Galactic center, the molecular gas exhibits distinctive properties than gas in the disk, and can be excited to high , and occasionally, very-high ratios (Oka et al., 1996, 1998b; Sawada et al., 2001). Overall, the ratio varies from 0.96 in the central kiloparsec region of the MW, while the typical disk value is 0.6-0.7 (Sawada et al., 2001). VHRG with is relatively rare and is typically found in the close vicinity of star formation, e.g. due to interaction with supernovae, or in the galactic center (Oka et al., 1996; Seta et al., 1998).
@@
There have been many efforts to characterize in external galaxies. However, previous measurements were often hindered by challenges in calibration, especially since intrinsically has a relatively small dynamic range (only a factor of 2-3 variation; Garcia-Burillo et al., 1993; Crosthwaite et al., 2002; Lundgren et al., 2004; Crosthwaite & Turner, 2007; Koda et al., 2012; Yajima et al., 2021; Leroy et al., 2022). The calibration issue often led to high over large regions of the galaxies, seemingly indicating widespread optically-thin CO gas, which is not very likely. Only HRG 0.9 in galactic centers has been a consistent and physically-plausible result (e.g., Braine & Combes, 1992; Leroy et al., 2009). Spatial variations in in external galaxies began to become apparent with improvements in calibration and techniques (e.g., Koda et al., 2012). In particular, with the advent of ALMA, studies of have become increasingly possible (Koda et al., 2020; Egusa et al., 2022; Maeda et al., 2022; den Brok et al., 2023). There have also been galaxy-wide studies of in normal galaxies, (e.g., Saintonge et al., 2017; Keenan et al., 2024, 2025), as well as in extreme systems such as starburst and (ultra)-luminous infrared galaxies (e.g., Papadopoulos et al., 2012; Montoya Arroyave et al., 2023).
@@
High-resolution studies of ( 100 pc) remain relatively scarce and are typically case-studies, or due to limited coverage, focus on particular regions in galaxies (Druard et al., 2014; Egusa et al., 2022; Maeda et al., 2022; den Brok et al., 2023; Koda et al., 2025). These studies have found values similar to those of Galactic studies as well as indications of trends with respect to galactic structures. Beyond case-studies, it is clear that a survey of in nearby galaxies at high-resolution and with well-calibrated data, is needed to understand the general evolution of molecular gas and their physical conditions. While there are existing surveys of variations (e.g., Yajima et al., 2021; Leroy et al., 2022; den Brok et al., 2025), they are at lower sensitivity and resolution, or the compared CO(1-0) and CO(2-1) data were often observed from different telescopes. As a result, overall systematic trends were not clear.
@@
In this paper, we analyze in twelve nearby galaxies using new high-resolution CO(1-0) data from the ALMA-FACTS survey and reimaged CO(2-1) data from the PHANGS-ALMA survey (Leroy et al., 2021). We present early results, which already empirically show systematic trends as a function of galactic structure. This is the third paper of the ALMA-FACTS Survey. on kpc-scales in the same sample of galaxies using the ALMA Total Power (TP) component has been studied (ALMA-FACTS Paper II; Komugi et al., 2025). We plan multi-wavelength comparisons for future studies.
II Observations and Data Reduction
@@
This section discusses the data reduction, mainly on the CO(2-1) data (Section II.1). The CO(1-0) data reduction is discussed in a separate paper (Koda et al. in prep.), but is practically the same with one note in Section II.2. This method has been discussed in depth and used by Koda et al. (2023) and Koda et al. (2025).
@@
Figure 1 shows the sample in Spitzer 3.6 m dust emission from S4G (Sheth et al., 2010; S4G Team, 2020) and SINGS (for NGC 3621, Kennicutt et al., 2003; SINGS Team, 2020) to demonstrate the field-of-view (FoV) of the observations. The FoVs of the 12m+07m CO(2-1) observations (yellow contours), and consequently the maps (section IV.1), cover the inner disks of the galaxies.
II.1 CO(2-1)
@@ II.1.1 Calibration
II.1.2 Imaging
The 12m and 07m array data for each galaxy were imaged together in MIRIAD, instead of CASA, because it takes into account spatial variations in the point spread function (PSF) (Koda et al., 2023). To prepare for imaging in MIRIAD, we used the CASA tasks split and mstransform to split the narrowband spectral window (SPW) containing the CO(2-1) emission out from each 12m and 07m MS, and then to regrid them to a common velocity grid with channel width 5 km/s. These were converted into MIRIAD format. We used the average system temperature Tsys and integration times tint from each MS following Koda et al. (2023). The invert task in MIRIAD was then performed on these 12m and 07m visibilities to produce the CO(2-1) data cubes. We used robust = +2 (approximately) corresponding to natural weighting), cellsize = 0.25”, and a velocity coverage and width of matching those of the CO(1-0) data cubes (Koda et al. in prep).
@@ II.1.2 Imaging
II.2 CO(1-0)
The data reduction for CO(1-0) is discussed in a separate paper (Koda et al., in prep.), but is practically the same as the one discussed above. The only difference is the very first step of imaging, where we did not have a reference cube to make a mask. We went through the full imaging process without any mask and generated a 12m+07m+TP cube in CO(1-0). A mask is then made with this cube for the final imaging which was done in the same way as discussed for CO(2-1).
III Maps
@@
To make the integrated intensity maps for each galaxy, we smoothed the CO(1-0) and CO(2-1) cubes to a common angular resolution of 2.5. We then regridded the CO(2-1) data spatially to the CO(1-0) data and applied a mask to the data cubes. The mask was based on the (cleaned) CO(1-0) signal-to-noise (SN) cube starting from volumes that have at least 80 pixels with 5.0. This mask was then expanded spatially by 2.5 (=10 pixels in length) to include potential diffuse extended emission (if it exists). The mask was not expanded in velocity. After applying this mask, the cubes were integrated along the velocity axis to produce the integrated intensity maps, and corresponding error maps were generated. The error in each pixel is calculated as RMS, where RMS is the RMS noise per channel width , and is the number of channels included in the velocity integration.
These are shown in Figures 3 and 4. The colorbar scales for are adjusted to cover 0.7 times the shown range of for a fairer comparison using a (fiducial) average line ratio of 0.7. The and maps appear similar for prominent structures, although their ratio maps show variations (see Section IV).
@@
(7-9) Adopted semi-major axis, position-angle, and ellipticity of the bar from Herrera-Endoqui et al. (2015) for bar region masks (section V.2.5).
@@
IV Results
From the integrated intensity maps in Section III, we calculated the line ratio maps (Figures 5-10). We applied a signal-to-noise (SN) cut of 5 in both CO(1-0) and CO(2-1). We summarize the values and their distribution for each galaxy individually. The spatial resolutions of our maps at the adopted angular resolution of 2.5 range from 50-200 pc amongst the sample (Table 1), larger than the typical size of a molecular cloud. The values presented here thus represent the average within individual molecular clouds.
IV.1 Maps and Distributions
The maps for each galaxy and their distributions (histograms), as well as their optical images for comparison, are shown in Figures 5-10. We summarize the variations in each galaxy separately (see Table 3). We tend to find similar trends in each morphological type and galactic structure (e.g., barred, unbarred, spiral), which leads us to classify them according to their optical and molecular gas morphologies (Section V).
We divide the molecular gas using the definitions adopted in Hasegawa (1997). Low-ratio gas (LRG) have (blue), high-ratio gas (HRG) are in between (red), and very-high-ratio gas (VHRG) have . The division of molecular gas at (white) is motivated both theoretically via the large-velocity gradient (LVG) approximation (Goldreich & Kwan, 1974; Scoville & Solomon, 1974) and observationally first in MW studies (e.g., Sakamoto et al., 1997; Oka et al., 1998a; Hasegawa, 1997). Typical molecular gas with physical conditions () = (300 cm-3, 10 K) have 0.6-0.7 (around the LRG-HRG boundary), with the transition from LRG to HRG requiring 2-3x denser and/or hotter gas (e.g., Koda et al., 2012). On the other hand, VHRG with is difficult to achieve, requiring optically thin conditions yet even denser/hotter gas.
We note that the shift from LRG to HRG (or vice-versa) is continuous in the values of , (and ) (Koda et al., 2025) and cannot be characterized by single numbers. Nevertheless, a crude guideline would be helpful to gauge the relative physical conditions of LRG and HRG. According to the calculations by Koda et al. (2012) which considered a range of , HRG requires 2-3 times higher and than LRG.
@@ IV.1 Maps and Distributions
IV.1.1 NGC 0628
NGC 0628 is an isolated spiral galaxy (SAc) with many star-forming regions (HII) along its spiral arms. It does not have a bar structure. At its distance, the angular resolution 2.5 corresponds to a spatial resolution of 119 pc. From Figure 3, the molecular gas distribution in NGC 0628 is relatively dim (e.g. compared to the other unbarred galaxies in the sample) in both transitions of CO and its spiral arms appear loosely wound. The stellar spiral arms extend beyond the CO coverage (e.g., Figure 1).
The measurements are concentrated around the spiral arms due to the sensitivity limitation, with that in the interarm regions only occasionally determined. The disk is dominated by low-ratio gas (Figure 5), encompassing 67% of the significantly detected pixels, while high-ratio gas makes up 26% (blue histogram). This is similar for the CO(1-0) intensity weighted distribution (green histogram), with low-ratio gas and high-ratio gas containing 77% and 19% respectively. Over the whole area, the mean is 0.65 while the intensity-weighted mean is 0.60.
IV.1.2 NGC 1097
@@
is high in the center, becomes lower along the bar, and then increases at the bar ends (Figure 5). Low-ratio gas and high-ratio gas make up 70% and 26% of the area (blue histogram), respectively. However, the CO(1-0) intensity contained by low-ratio and high-ratio gas are similar, at 47% and 50% respectively (green histogram). Over the whole area, the mean is 0.63 while the intensity-weighted mean is 0.71.
@@ IV.1.3 NGC 1512
is high in the center and becomes lower along the offset ridges (Figure 6). The along the spiral arms fluctuate between low-ratio and high-ratio, with some high-ratio gas spatially coincident with HII regions. This galaxy also has a central gas concentration and high ratio region larger than the nominal definition of the central 1 kpc. Indeed, the histogram for the central 2 kpc explains most of the HRG. Low-ratio and high-ratio gas encompass 50% and 42% of the area shown, respectively. However, low-ratio gas contain 31% of the CO(1-0) intensity, while 61% is contained within the high-ratio gas. Over the whole area, the mean is 0.72 while the intensity-weighted mean is 0.79. From the histograms, the central 1-2 kpc ( kpc) contain mostly high-ratio gas.
IV.1.4 NGC 3351
@@
In Figure 6, is high in the center and decreases along the offset ridges, but fluctuates in the surrounding spiral arms between low and high ratios. Low-ratio and high-ratio gas make up 58% and 38% of the area shown (blue histogram). However, the amount of low-ratio gas and high-ratio gas contained by the CO(1-0) intensity are 36% and 62%, respectively (green histogram). Over the whole area, the mean is 0.69 while the intensity-weighted mean is 0.77. From the histograms, high-ratio gas is concentrated predominantly in the central kiloparsec ( kpc) region.
Teng et al. (2022) studied , and line ratios (at 2.1 resolution; 100 pc) in the central kiloparsec of NGC 3351. They also found that the central region consists of high-ratio gas .
IV.1.5 NGC 3521